respected sir,
1] A store’s selling price of $2,240 for a certain computer would yield a profit of 40 percent of the store’s cost for the computer. What selling price would yield a profit of 50 percent of the computer’s cost?

plz solve my math sum...
thank u .

Answer by RAY100(1637)

Let x = cost
.
1.4 * x = $ 2240
.
x= $ 1600
.
1.5 * x = 1.5 * 1600 = $ 2400
.
check
.
40 % of 1600 is ,,,, (.4) *1600 = 640
640 + 1600 = $ 2240,,,,,ok
,
50 % of 1600 = 800
800 + 1600 = $ 2400,,,,ok